OpenStudy (anonymous):

Solve the system of equations: 2x-3y=-12 and x-y=4. Please help

OpenStudy (anonymous):

you can use the substitution method for these. Solve an equation (either one is fine) for y, so that the equation is in the y = mx + b format.

OpenStudy (anonymous):

When you solve the equation, plug the "mx + b" part into the separate equation that you haven't manipulated yet. So, if at first you use "x - y = 4", use "2x - 3y = -12" at this point.

OpenStudy (anonymous):

Solve the "2x - 3y = -12" for x, after you insert the "y = mx + b" equation

OpenStudy (anonymous):

Got it thank you.